OpenStudy (anonymous):

What is the area 16m of a circle use 3.14 to get the answer?

OpenStudy (anonymous):

is 16m the diameter?

OpenStudy (anonymous):

It just says 16m

OpenStudy (shadowlegendx):

Do you know the formula for finding the area of a circle?

OpenStudy (anonymous):

No lol math is not my thing. I hate math I am slow.

OpenStudy (shadowlegendx):

\[A = \pi \times r^{2} \]

OpenStudy (shadowlegendx):

A = Area \[\pi = 3.14 \] r = radius

OpenStudy (shadowlegendx):

The confusing part of this question is we do not know if 16m is the radius or diameter. In your original question can you see any of those word phrases?

OpenStudy (shadowlegendx):

@katebored are you there?

OpenStudy (anonymous):

There is a picture of a circle next to it is a outside line that says 16m next to it.

OpenStudy (anonymous):

Think its the height